(x+a)^8=x^8(1+a/x)^8
Pls describe.....
1
Expert's answer
2019-05-30T11:12:32-0400

(x + a)8 = x8 + 8x7a + 28x6a2 + 56x5a3 + 70x4a4 + 56x3a5 + 28x2a6 + 8xa7 + a8


(1 + a/x)8 = 1 + 8(a/x) + 28(a/x)2 + 56(a/x)3 + 70(a/x)4 + 56(a/x)5 + 28(a/x)6 + 8(a/x)7 + (a/x)8


x8(1 + a/x)8 = x8 + 8ax8/x + 28a2x8/x2 + 56a3x8/x3 + 70a4x8/x4 + 56a5x8/x5 + 28a6x8/x6 + 8a7x8/x7 + a8x8/x8 = x8 + 8x7a + 28x6a2 + 56x5a3 + 70x4a4 + 56x3a5 + 28x2a6 + 8xa7 + a8 => (x + a)8 = x8(1 + a/x)8.